OpenStudy (anonymous):

Another sequence. The fractions are throwing me off 1,-1/2,1/4

OpenStudy (trantom):

This is a geometric sequence. Do you know why?

OpenStudy (anonymous):

I forgot

OpenStudy (trantom):

A geometric sequence is a sequence of numbers you can use a common factor (something you multiply by) to get the next number. Lets start with what you multiply 1 by to get -1/2. Do you know?

OpenStudy (trantom):

Well, if you don't, think of it like this. Dividing by 2 is the same as multiplying by 1/2. 1 times 1/2 = 1/2. Now make that negative. 1*-1/2=-1/2

OpenStudy (trantom):

Also, a negative times a negative equals a positive so -1/2*-1/2= +1/4

OpenStudy (trantom):

So the recursive formula for this sequence is A(n)=A(n-1)(-1/2)

OpenStudy (trantom):

and A(1)= 1

OpenStudy (usukidoll):

I think a faster way of getting the key formula would be (-1)^n 1/2^n

OpenStudy (usukidoll):

so if we let n = 1 (-1)^1 1/2^1 then it's -1/2 n = 2 (-1)^2 1/2^2 = 1/4 and so on and so on

OpenStudy (usukidoll):

the pattern is positive negative positive negative... so we start at n = 0

OpenStudy (usukidoll):

you need this === > (-1)^n 1/2^n and let n = 0., 1, 2, ....

OpenStudy (anonymous):

Ah ok